### v1

@@ 1 Introduction
Many such anecdotal examples could be recounted and the literature reports numerous studies on the problem-solving skills of LLMs [chen2024, huang2024, jiang2023, leng2025]. OpenAI researchers likely also had a noosemic experience when they learned about the emergent features of GPT-3 [wei2022emergent], which, with a neural architecture similar to that of GPT-2 but trained on a much larger dataset, showed such emergent features that they declared in the paper that “Language Models Are Few-Shot Learners” [brown2020language]. The title highlights that this emergent behavior was not merely an incremental improvement, but a surprising discovery. Before GPT-3, the dominant paradigm was pre-training followed by extensive fine-tuning. The idea that a model could perform new and complex tasks with only a few examples (few-shot), or even none (zero-shot), provided directly in the prompt was revolutionary. The paper’s structure itself, focused on exploring these unexpected abilities across dozens of benchmarks rather than on architectural novelty, further underscores this sense of discovery [^10] In an interview, when asked about the model’s emergent abilities, OpenAI CEO Sam Altman confirmed this sense of wonder, stating: “[…] for some definition of reasoning […] it can do some kind of reasoning. And I think that’s remarkable and the thing that’s most exciting and somehow out of ingesting human knowledge, it’s coming up with this reasoning capability…” [Fridman2023Altman]. See https://www.youtube.com/watch?v=L_Guz73e6fw.. Nevertheless, it seems that the initial “wow effect” underlying the noosemic experience is almost always followed by an “effect of habituation”. As users become accustomed to the system’s level of emulated intelligence, the surprise fades. The “wow effect” is further diminished whenever the user encounters a cognitive limitation of the model—when it repeatedly fails to understand a request or generates repetitive, inaccurate text, or in general “hallucinates”. Such situations induce a sense of frustration that belongs to the semantic field of noosemia in its negative sense. In other words, the “effect of habituation” pertains to the semantic domain of the a-noosemic experience. Nevertheless, given the ongoing, substantial improvement of generative AI systems and the advent of agentic architectures, it is reasonable to expect that the near future will be marked by alternating episodes of noosemic surprise and a-noosemic familiarity. While this study focuses on the noosemic experience, it will also provide a clear definition of a-noosemia.
@@
Ultimately, this study connects the human experience of generative artificial intelligence with its technical constitution, exploring the explanatory gap wherein symbolic ambivalence and the negotiation of meaning reside. The ongoing cognitive revolution demonstrates how humanity, faced with a “forest of signs and symbols” born from the objectification of knowledge and the bureaucratization of the world, is now creating a technology to navigate it. In these terms, the AI of the future can be understood as an enhancement of the noosphere [teilhard1959phenomenon], destined to modify technology as we perceive it today. While machines have limitations, human technique (technē) [^11] We distinguish between technique technē, the artful practice of human original creation and the resulting technology, which embodies it and is an objectification of technique. is precisely the practice capable of overcoming them—a pattern that history consistently demonstrates.
@@
The remainder of the paper is structured as follows: Section 2 introduces and formalizes the concept of Noosemia, detailing its mechanisms, manifestations, etymological foundations, and contrasting it with the phenomenon of a-noosemia. A discussion on the possible cues causing noosemia in the structure of Transformer architectures and its components is provided. Section LABEL:sec:background provides an overview of the main theories and empirical findings on the attribution of mind and meaning in human–AI interaction, including a reappraisal of the Turing Test, the intentional stance, and recent studies on Theory of Mind in language models. Section LABEL:sec:gap investigates the contemporary explanatory gap, focusing on its semiotic, symbolic and systemic dimensions, and explores how complexity and opacity foster new forms of cognitive projection and sense-making. Section LABEL:sec:comparisons compares Noosemia to related cognitive phenomena such as pareidolia, animism and the uncanny valley, elucidating its unique characteristics. Finally, in Section LABEL:sec:From_Digital_Enaction_to_Embodied_Mind we will try to take a look at the future by following the major developments underway and the dedicated research lines also in relation to a convergence between the proposed architectures and cognitive models. The research concludes by outlining future research directions and the broader theoretical and practical implications of Noosemia for human–AI interaction.
@@ 2 Noosemia: Definition and Theoretical Foundation
Noosemic Phenomenon: Mechanisms and Manifestations
@@
To address this conceptual gap, we introduce the notion of Noosemia [^12] From the Greek nous (mind, intellect) and semìa (sign, mark)., a new term describing the specific projection of mind, intentionality, or interiority onto generative AI systems on the basis of their semantic fluency, dialogic creativity, and epistemic opacity. Unlike classical anthropomorphism, Noosemia is not triggered by physical or perceptual features, but emerges from the capacity of these systems to generate coherent, context-sensitive, and often unexpectedly meaningful language, thus synchronizing with and sometimes anticipating the cognitive patterns of the human interlocutor. Therefore, a paradigmatic instance of Noosemia can occur during a first encounter with an advanced conversational agent. For example, a user may pose a complex or ambiguously phrased question to a generative AI and receive a response that resolves the ambiguity and demonstrates creative inference or analogical reasoning that aligns with, or even surpasses, the user’s expectations. It is assumed that the immediate reaction—often one of astonishment or disorientation—reflects the sudden attribution of interiority to the system, despite the user’s awareness of its artificial nature. Such moments, characterized by an “wow effect” (see Section 1), exemplify the distinctive projection of mind that Noosemia aims to describe.
@@
Definition: Noosemia Noosemìa (from Greek noûs – mind, and sēmeîon – sign) is a cognitive and phenomenological phenomenon in which a human, interacting with a generative artificial intelligence system—especially those capable of dialogic or multimodal interaction—typically attributes mental states, intentionality, and a sense of interiority (i.e., inwardness or subjective presence) to the system. Distinct from classical animism, Noosemia inherits a cognitive disposition toward agency projection, now elicited by linguistic performance rather than physical appearance. This projection arises from the epistemic opacity and perceived “magical” quality of the AI’s outputs, often resulting in moments of genuine cognitive astonishment. Noosemìa thus emerges within a domain of symbolic ambivalence, where the boundary between sign and mind becomes blurred and meaning is co-constructed in a dialogic and open-ended manner. The phenomenon is technically grounded in the emergent properties—such as hierarchy, complexity, and opacity—of advanced generative architectures, which enable such projection. Noosemìa refers to a phenomenological attribution by the user, not to the actual presence of consciousness or intentionality in the artificial system.
@@
This new term is motivated by the need of a suited specification, within the generative AI era, of established frameworks—such as Dennett’s intentional stance [dennett1987intentional] and the inadequacy of classical animism [mauss1906origins]—to fully account for the emergent projection of agency observed in contemporary interactions with LLMs. Traditional anthropomorphism is largely based on visual or behavioral cues, while Noosemìa is fundamentally linguistic and dialogic and can cover a wide spectrum of expressive forms, which machines are and will be able to master (such as multimodal interaction, agency emulation, etc.). In this context, the machine is experienced not as a mechanical automaton, but as an interlocutor capable of producing meaning, forming analogies and even surprising the user with novel semantic connections and forms of creativity [desantis2023apocalissi]. Such semantic connections may occur at the level of the semantic field carried by words within the generative flow of the LLM. More rarely, they can also emerge at the level of neologism creation, a phenomenon enabled by the very nature of the basic semantic units employed by LLMs—namely, tokens, which are subwords or fragments of words. This phenomenon has been empirically documented by both [kaplan2024innerlexicon] and [malkin2021gpt], which show how LLMs can generate and interpret neologisms through the composition of subword tokens, often producing plausible meanings even for words never seen during training (Out-Of-Vocabulary-Words).
@@
This projection is catalyzed by the intrinsic opacity of modern AI systems, whose internal processes remain inaccessible even to most experts — as for our brain [chalmers1996conscious], thus fostering an explanatory gap between observable linguistic competence and the nontransparent causal substrate. As a result, users often experience a sense of cognitive resonance (“wow effect” [^13] For a discussion on the “ELIZA effect” see Section 1.), a phenomenological shock that could prompt the attribution of proto-mind to the system. More generally, Noosemìa refers to the experience in which, during a dialogue with an AI (or a more general interaction), the user detects a spark of intelligence [bubeck2023]—an impression that leaves them momentarily amazed and perplexed. This is akin to the cognitive dynamics that underpin the art of prestidigitation.
@@
Noosemìa thus designates a new interpretive horizon for human–generative AI interaction, one in which meaning, understanding and agency are co-constructed in the interplay of linguistic performance and epistemic uncertainty. It provides both a descriptive and an analytical tool for exploring the cognitive, ethical, systemic and epistemological consequences of living with systems whose outputs are intelligible but whose workings are not. In fact, it is well-known that even many AI researchers and practitioners acknowledge that, despite designing and training these systems, they often lack the tools or frameworks to fully explain why a LLM generates a specific output. This is not merely a practical issue, but an inherent feature of overparameterized, non-linear architecture, which is in all respects a complex system and as such shows typical characteristics such as emergency [wei2022emergent, chen2024quantifying, teehan2022emergent, manning2020emergent]. As noted in a comprehensive review [li2021interpretable], the inner decision-making processes of deep neural networks frequently remain opaque even to those who built them.
@@
A full exploration of the historical, psychological, and sociocultural roots of mind attribution, as well as comparisons with related phenomena such as animism, pareidolia, and the uncanny valley, are developed in the following sections. Here, our primary aim is to formally introduce Noosemìa as a necessary concept for capturing the specific mode of sense-making and agency projection elicited by advanced generative AI – see Fig. 2 for a pictorial representation.
@@
Figure 2: Conceptual schema of Noosemìa. In a typical interaction, the user provides a linguistic input. The generative AI system (LLM) produces a fluent and context-sensitive output, which passes through a zone of epistemic opacity (“black box” or symbolic ambivalence). This opacity triggers the cognitive projection of mind, agency, or even interiority onto the system—the essence of Noosemìa. The process is dialogic and iterative, evolving with each exchange.
A-Noosemia: Absence and Collapse of Noosemic Projection
@@
Noosemìa (Attribution of mind/agency) Repeated Errors / Mechanical Failures Loss of “Wow” Effect / Familiarity Epistemic Skepticism / Disbelief Excessive Opacity / Black-Boxing A-Noosemìa (Suspension of attribution)
Figure 3: Main cognitive and technical pathways leading from Noosemìa (attribution of mind/agency) to A-Noosemìa (suspension or collapse of attribution) in human–AI interaction.
It is important to note that a-noosemia, as conceptualized here, is not merely the absence of anthropomorphic projection, but the collapse of a more subtle, semiotic and phenomenological engagement. The user is no longer caught in the spiral of fascination and meaning-making; instead, they become immune to the system’s cues and disengage from the dialogic construction of agency. In this sense, a-noosemia completes the dialectic initiated by noosemia, revealing the contingent, fragile, and ultimately reversible nature of the attribution of mind in human–AI interaction. Moreover, the analytic contrast between noosemia and a-noosemia enriches our understanding of the dynamics of sense-making in the age of generative AI, drawing attention to the conditions under which the boundaries between tool and interlocutor are alternately dissolved and reasserted. In empirical and practical terms, the prevalence of a-noosemic states may help to explain phenomena such as user disengagement, loss of trust, or the emergence of critical or skeptical attitudes towards artificial systems, especially as the novelty of interaction fades and the complexity of tasks increases. To illustrate these distinctions more concretely, Table 1 contrasts the core features of noosemia and a-noosemia, highlighting how these opposing experiences shape the dynamics of sense-making and agency attribution in contemporary human–AI interactions.
@@
Etymological and Conceptual Foundations of the Term “Noosemìa”
@@
To fully appreciate the specific scope of Noosemìa141414 The term Noosemìa retains the grave accent on the “ì” to reflect its etymological derivation from the Greek and to guide proper pronunciation. However, in English-language contexts and for indexing purposes, the unaccented form Noosemia is equally acceptable and may be used interchangeably throughout the text., it is essential to analyze the historical and semantic origins of its components. The construction of the neologism is grounded in a deliberate synthesis of two foundational Greek morphemes, each bearing significant weight in the history of philosophy and semiotics. The root noûs ($\nu o\acute{\upsilon}\varsigma$) refers to the mind, intellect, or the principle of reason—a concept that, since Anaxagoras, has denoted the capacity for ordering, discerning, and understanding the world [anaxagoras1982fragments, aristotle1984metaphysics]. For Aristotle, noûs represented the highest faculty of the mind, that which enables humans to apprehend universal truths and meanings. In subsequent philosophical traditions, from Neoplatonism to medieval scholasticism, noûs was associated with both divine intellect and the human capacity for abstraction and interiority. The second morpheme, the suffix -semìa, is derived from the Greek word sēmeîon ($\sigma\eta\mu\epsilon\acute{\iota}o\nu$) [^15] The suffix ‘-semìa’ is derived from the Greek sēmeîon ($\sigma\eta\mu\epsilon\acute{\iota}o\nu$) in deliberate contrast to sêma ($\sigma\widehat{\eta}\mu\alpha$). While both terms can be translated as ‘sign,’ sēmeîon specifically denotes a sign that is relational and requires interpretation; it functions as an indicator or trace pointing to an inferred, underlying reality. Conversely, sêma refers to a more concrete, self-contained sign, such as a physical marker or a monument, where the signifier and signified are more closely identified. This distinction is fundamental, as Noosemia describes the process of interpreting an AI’s outputs as sēmeîa—signs that prompt the user to infer the existence of an unobserved mind—rather than perceiving them as a direct embodiment of that mind (a sêma). and designates a sign, signal, or mark—a central notion in both ancient rhetoric and the theory of signs. In the work of Hippocrates and later Galen, $\sigma\eta\mu\epsilon\acute{\iota}o\nu$ was used to describe signs and symptoms, while in the field of semiotics, the concept became central to the understanding of how meaning is produced, transmitted, and interpreted [eco1976theory]. Modern semiotic theory, notably that of Charles Peirce and Umberto Eco, situates the sign as the foundational unit of meaning-making, mediating between the world, the mind, and language [peirce1931collected, eco1976theory]. By combining noûs and sēmeîon, the neologism Noosemìa encapsulates the projection or emergence of “mind” as it arises specifically through the mediation of signs. This etymological choice underscores the phenomenon’s fundamentally cognitive and semiotic dimensions. In fact, Noosemìa does not denote an attribution of mind in a vacuum, but rather as a response to the complex, sign-generating capacities of generative AI systems. In this sense, the term aligns with a lineage of philosophical inquiry into how intellect and meaning are always-already entangled with signs and symbolic structures [eco1976theory, aristotle1984metaphysics, DeSantis2021].
Philosophical and Systemic Foundations

### v2

@@ 1 Introduction
Many such anecdotal examples could be recounted and the literature reports numerous studies on the problem-solving skills of LLMs [190, 237, 241, 254]. OpenAI researchers likely also had a noosemic experience when they learned about the emergent features of GPT-3 [329], which, with a neural architecture similar to that of GPT-2 but trained on a much larger dataset, showed such emergent features that they declared in the paper that “Language Models Are Few-Shot Learners” [182]. The title highlights that this emergent behavior was not merely an incremental improvement, but a surprising discovery. Before GPT-3, the dominant paradigm was pre-training followed by extensive fine-tuning. The idea that a model could perform new and complex tasks with only a few examples (few-shot), or even none (zero-shot), provided directly in the prompt was revolutionary. The paper’s structure itself, focused on exploring these unexpected abilities across dozens of benchmarks rather than on architectural novelty, further underscores this sense of discovery [^10] In an interview, when asked about the model’s emergent abilities, OpenAI CEO Sam Altman confirmed this sense of wonder, stating: “[…] for some definition of reasoning […] it can do some kind of reasoning. And I think that’s remarkable and the thing that’s most exciting and somehow out of ingesting human knowledge, it’s coming up with this reasoning capability…” [219]. See https://www.youtube.com/watch?v=L_Guz73e6fw.. Nevertheless, it seems that the initial “wow effect” underlying the noosemic experience is almost always followed by an “effect of habituation”. As users become accustomed to the system’s level of emulated intelligence, the surprise fades. The “wow effect” is further diminished whenever the user encounters a cognitive limitation of the model—when it repeatedly fails to understand a request or generates repetitive, inaccurate text, or in general “hallucinates”. Such situations induce a sense of frustration that belongs to the semantic field of noosemia in its negative sense. In other words, the “effect of habituation” pertains to the semantic domain of the a-noosemic experience. Nevertheless, given the ongoing, substantial improvement of generative AI systems and the advent of agentic architectures, it is reasonable to expect that the near future will be marked by alternating episodes of noosemic surprise and a-noosemic familiarity. While this study focuses on the noosemic experience, it will also provide a clear definition of a-noosemia.
@@
Ultimately, this study connects the human experience of generative artificial intelligence with its technical constitution, exploring the explanatory gap wherein symbolic ambivalence and the negotiation of meaning reside. The ongoing cognitive revolution demonstrates how humanity, faced with a “forest of signs and symbols” born from the objectification of knowledge and the bureaucratization of the world, is now creating a technology to navigate it. In these terms, the AI of the future can be understood as an enhancement of the noosphere [313], destined to modify technology as we perceive it today. While machines have limitations, human technique (technē) [^11] We distinguish between technique (technē), the artful practice of human original creation and the resulting technology, which embodies it and is an objectification of technique. is precisely the practice capable of overcoming them—a pattern that history consistently demonstrates.
@@
The remainder of the paper is structured as follows. Section 2 introduces and formalizes the concept of Noosemia, detailing its mechanisms, manifestations, etymological foundations, and contrasting it with the phenomenon of a-noosemia. A discussion on the possible cues causing noosemia in the structure of Transformer architectures and its components is provided. Section 3 provides an overview of the main theories and empirical findings on the attribution of mind and meaning in human–AI interaction, including a reappraisal of the Turing Test, the intentional stance, and recent studies on Theory of Mind in language models. Section 4 investigates the contemporary explanatory gap, focusing on its semiotic, symbolic and systemic dimensions, and explores how complexity and opacity foster new forms of cognitive projection and sense-making. Section 5 compares Noosemia to related cognitive phenomena such as pareidolia, animism and the uncanny valley, elucidating its unique characteristics. Finally, in Section 6 we will try to take a look at the future by following the major developments underway and the dedicated research lines also in relation to a convergence between the proposed architectures and cognitive models. The research concludes by outlining future research directions and the broader theoretical and practical implications of Noosemia for human–AI interaction.
@@ 2 Noosemia: Definition and Theoretical Foundation
2.1 Noosemic Phenomenon: Mechanisms and Manifestations
@@
To address this conceptual gap, we introduce the notion of Noosemia [^12] From the Greek noûs (mind, intellect) and sēmeîon (sign, mark)—see Section 2.3., a new term describing the specific projection of mind, intentionality, or interiority onto generative AI systems on the basis of their semantic fluency, dialogic creativity, and epistemic opacity. Unlike classical anthropomorphism, Noosemia is not triggered by physical or perceptual features, but emerges from the capacity of these systems to generate coherent, context-sensitive, and often unexpectedly meaningful language, thus synchronizing with and sometimes anticipating the cognitive patterns of the human interlocutor. Therefore, a paradigmatic instance of Noosemia can occur during a first encounter with an advanced conversational agent. For example, a user may pose a complex or ambiguously phrased question to a generative AI and receive a response that resolves the ambiguity and demonstrates creative inference or analogical reasoning that aligns with, or even surpasses, the user’s expectations. It is assumed that the immediate reaction—often one of astonishment or disorientation—reflects the sudden attribution of interiority to the system, despite the user’s awareness of its artificial nature. Such moments, characterized by a “wow effect” (see Section 1), exemplify the distinctive projection of mind that Noosemia aims to describe.
@@
Definition: Noosemia Noosemìa (from Greek noûs – mind, and sēmeîon – sign) is a cognitive and phenomenological pattern in which a human, interacting with a generative artificial intelligence system—especially those capable of dialogic or multimodal interaction—typically attributes mental states, intentionality, and a sense of interiority (i.e., inwardness or subjective presence) to the system. Distinct from classical animism, Noosemia inherits a cognitive disposition toward agency projection, now elicited by linguistic performance rather than physical appearance. This projection arises from the epistemic opacity and perceived “magical” quality of the AI’s outputs, often resulting in moments of genuine cognitive astonishment. Noosemia thus emerges within a domain of symbolic ambivalence, where the boundary between sign and mind becomes blurred and meaning is co-constructed in a dialogic and open-ended manner. The phenomenon is technically grounded in the emergent properties—such as hierarchy, complexity, and opacity—of advanced generative architectures, which enable such projection. Noosemia refers to a phenomenological attribution by the user, not to the actual presence of consciousness or intentionality in the artificial system.
@@
This new term is motivated by the need of a suited specification, within the generative AI era, of established frameworks—such as Dennett’s intentional stance [207] and the inadequacy of classical animism [271]—to fully account for the emergent projection of agency observed in contemporary interactions with LLMs. Traditional anthropomorphism is largely based on visual or behavioral cues, while Noosemia is fundamentally linguistic and dialogic and can cover a wide spectrum of expressive forms, which machines are and will be able to master (such as multimodal interaction, agency emulation, etc.). In this context, the machine is experienced not as a mechanical automaton, but as an interlocutor capable of producing meaning, forming analogies and even surprising the user with novel semantic connections and forms of creativity [203]. Such semantic connections may occur at the level of the semantic field carried by words within the generative flow of the LLM. More rarely, they can also emerge at the level of neologism creation, a phenomenon enabled by the very nature of the basic semantic units employed by LLMs—namely, tokens, which are subwords or fragments of words. This phenomenon has been empirically documented by both [245] and [263], which show how LLMs can generate and interpret neologisms through the composition of subword tokens, often producing plausible meanings even for words never seen during training (Out-Of-Vocabulary-Words).
@@
This projection is catalyzed by the intrinsic opacity of modern AI systems, whose internal processes remain inaccessible even to most experts — as for our brain [187], thus fostering an explanatory gap between observable linguistic competence and the nontransparent causal substrate. As a result, users often experience a sense of cognitive resonance (“wow effect” [^13] For a discussion on the “ELIZA effect” see Section 1.), a phenomenological shock that could prompt the attribution of proto-mind to the system. More generally, Noosemia refers to the experience in which, during a dialogue with an AI (or a more general interaction), the user detects a spark of intelligence [183]—an impression that leaves them momentarily amazed and perplexed. This is akin to the cognitive dynamics that underpin the art of prestidigitation.
@@
Noosemia thus designates a new interpretive horizon for human–generative AI interaction, one in which meaning, understanding and agency are co-constructed in the interplay of linguistic performance and epistemic uncertainty. It provides both a descriptive and an analytical tool for exploring the cognitive, ethical, systemic and epistemological consequences of living with systems whose outputs are intelligible but whose workings are not. In fact, it is well-known that even many AI researchers and practitioners acknowledge that, despite designing and training these systems, they often lack the tools or frameworks to fully explain why a LLM generates a specific output. This is not merely a practical issue, but an inherent feature of overparameterized, non-linear architecture, which is in all respects a complex system and as such shows typical characteristics such as emergency [329, 188, 312, 264]. As noted in a comprehensive review [257], the inner decision-making processes of deep neural networks frequently remain opaque even to those who built them.
@@
A full exploration of the historical, psychological, and sociocultural roots of mind attribution, as well as comparisons with related phenomena such as animism, pareidolia, and the uncanny valley, are developed in the following sections. Here, our primary aim is to formally introduce Noosemia as a necessary concept for capturing the specific mode of sense-making and agency projection elicited by advanced generative AI – see Fig. 2 for a pictorial representation.
@@
Figure 2: Conceptual schema of Noosemia. In a typical interaction, the user provides a linguistic input. The generative AI system (LLM) produces a fluent and context-sensitive output, which passes through a zone of epistemic opacity (“black box” or symbolic ambivalence). This opacity triggers the cognitive projection of mind, agency, or even interiority onto the system—the essence of Noosemia. The process is dialogic and iterative, evolving with each exchange.
2.2 A-Noosemia: Absence and Collapse of Noosemic Projection
@@
Noosemia (Attribution of mind/agency) Repeated Errors / Mechanical Failures Loss of “Wow” Effect / Familiarity Epistemic Skepticism / Disbelief Excessive Opacity / Black-Boxing A-Noosemia (Suspension of attribution)
Figure 3: Main cognitive and technical pathways leading from Noosemia (attribution of mind/agency) to A-Noosemia (suspension or collapse of attribution) in human–AI interaction.
It is important to note that a-noosemia, as conceptualized here, is not merely the absence of anthropomorphic projection, but the collapse of a more subtle, semiotic and phenomenological engagement. The user is no longer caught in the spiral of fascination and meaning-making; instead, he becomes immune to the system’s cues and disengage from the dialogic construction of agency. In this sense, a-noosemia completes the dialectic initiated by noosemia, revealing the contingent, fragile, and ultimately reversible nature of the attribution of mind in human–AI interaction. Moreover, the analytic contrast between noosemia and a-noosemia enriches our understanding of the dynamics of sense-making in the age of generative AI, drawing attention to the conditions under which the boundaries between tool and interlocutor are alternately dissolved and reasserted. In empirical and practical terms, the prevalence of a-noosemic states may help to explain phenomena such as user disengagement, loss of trust, or the emergence of critical or skeptical attitudes towards artificial systems, especially as the novelty of interaction fades and the complexity of tasks increases. To illustrate these distinctions more concretely, Table 1 contrasts the core features of noosemia and a-noosemia, highlighting how these opposing experiences shape the dynamics of sense-making and agency attribution in contemporary human–AI interactions.
@@
2.3 Etymological and Conceptual Foundations of the Term “Noosemìa”
@@
To fully appreciate the specific scope of Noosemìa141414 The term Noosemìa retains the grave accent on the “ì” to reflect its etymological derivation from the Greek and to guide proper pronunciation. However, in English-language contexts and for indexing purposes, the unaccented form Noosemia is equally acceptable and may be used interchangeably throughout the text., it is essential to analyze the historical and semantic origins of its components. The construction of the neologism is grounded in a deliberate synthesis of two foundational Greek morphemes, each bearing significant weight in the history of philosophy and semiotics. The root noûs ($\nu o\acute{\upsilon}\varsigma$) refers to the mind, intellect, or the principle of reason—a concept that, since Anaxagoras, has denoted the capacity for ordering, discerning, and understanding the world [170, 172]. For Aristotle, noûs represented the highest faculty of the mind, that which enables humans to apprehend universal truths and meanings. In subsequent philosophical traditions, from Neoplatonism to medieval scholasticism, noûs was associated with both divine intellect and the human capacity for abstraction and interiority. The second morpheme, the suffix -semìa, is derived from the Greek word sēmeîon ($\sigma\eta\mu\epsilon\acute{\iota}o\nu$) [^15] The suffix ‘-semìa’ is derived from the Greek sēmeîon ($\sigma\eta\mu\epsilon\acute{\iota}o\nu$) in deliberate contrast to sêma ($\sigma\widehat{\eta}\mu\alpha$). While both terms can be translated as ‘sign,’ sēmeîon specifically denotes a sign that is relational and requires interpretation; it functions as an indicator or trace pointing to an inferred, underlying reality. Conversely, sêma refers to a more concrete, self-contained sign, such as a physical marker or a monument, where the signifier and signified are more closely identified. This distinction is fundamental, as Noosemia describes the process of interpreting an AI’s outputs as sēmeîa—signs that prompt the user to infer the existence of an unobserved mind—rather than perceiving them as a direct embodiment of that mind (a sêma). and designates a sign, signal, or mark—a central notion in both ancient rhetoric and the theory of signs. In the work of Hippocrates and later Galen, $\sigma\eta\mu\epsilon\acute{\iota}o\nu$ was used to describe signs and symptoms, while in the field of semiotics, the concept became central to the understanding of how meaning is produced, transmitted, and interpreted [210]. Modern semiotic theory, notably that of Charles Peirce and Umberto Eco, situates the sign as the foundational unit of meaning-making, mediating between the world, the mind, and language [291, 210]. By combining noûs and sēmeîon, the neologism Noosemia encapsulates the projection or emergence of “mind” as it arises specifically through the mediation of signs. This etymological choice underscores the phenomenon’s fundamentally cognitive and semiotic dimensions. In fact, Noosemia does not denote an attribution of mind in a vacuum, but rather as a response to the complex, sign-generating capacities of generative AI systems. In this sense, the term aligns with a lineage of philosophical inquiry into how intellect and meaning are always-already entangled with signs and symbolic structures [210, 172, 202].
2.4 Philosophical and Systemic Foundations